# Rotating black holes in a class of scalar-Gauss-Bonnet gravity

### v1

@@ Abstract
In this study, we investigate rotating black hole solutions within a scalar–Gauss–Bonnet gravity framework that incorporates a quadratic Gauss–Bonnet term. By employing a quadratic–exponential coupling function between the scalar field and the Gauss–Bonnet invariant, we derive both the standard General Relativity solutions and novel scalarized black hole configurations. Utilizing a pseudo-spectral method to solve the coupled field equations, we examine how black hole spin and coupling constants influence the existence and properties of these solutions. Our findings reveal that both the rotation of the black hole and the quadratic coupling term effectively constrain the parameter space available for scalarization. Moreover, we demonstrate that, over a wide range of parameters, scalarized black holes exhibit higher entropy than Kerr black holes of equivalent mass and spin, indicating that they are thermodynamically favored. These results significantly expand the phase space of black holes in modified gravity theories.
@@ I Introduction
Generally, scalarization is unattainable for BHs with intermediate and supermassive mass scales [31, 32, 33, 34, 35, 36]. Consquently, the prevailing expectation is that BHs in these mass ranges should be well described by the Kerr metric, even when solar-mass BHs are not. However, Ref. [63] considers a model in which the Gauss-Bonnet term is quadratically coupled to a scalar field. By studying the model in spherical symmetry, they found that unlike in the conventional framework, BHs can undergo spontaneous scalarization within a finite mass window, which may include supermassive BHs. Such deviations are particularly interesting because they offer a potential observation window through future gravitational wave detectors like LISA, Taiji, and TianQin to probe into the underlying gravitational theory [64, 65, 66, 67].
In the present paper, we extend these investigations to rotating BHs. By employing a quadratic-exponential coupling function, we explore how the inclusion of a quadratic Gauss-Bonnet term modifies the spectrum of rotating BH solutions. Our study addresses two central questions: How does the presence of angular momentum affect the onset and domain of scalarization? And how do quadratic curvature corrections influence the stability and thermodynamic properties of these scalarized configurations compared to their Kerr counterparts? To answer these questions, we numerically solve the full set of modified field equations under stationary and axisymmetric assumptions using the Chebyshev pseudo-spectral method combined with a Newton-Raphson iterative scheme. Our analysis reveals that both the spin parameter and the additional quadratic coupling act to suppress the scalarization, thereby reducing the parameter space in which scalarized solutions exist. Nevertheless, in many regimes the scalarized BHs exhibit higher entropy than Kerr BHs of the same mass and spin, suggesting that they are thermodynamically favored.
@@ II.1 Action and basic equations
We consider a four-dimensional sGB gravity theory with quadratic curvature corrections, defined by the action [63]:
@@
[…]
@@ II.2 Equations of motion and boundary behaviors
For numerical implementation of the Einstein-scalar system via the Chebyshev pseudo-spectral method, we introduce compactified radial coordinate $x=1-2r_{H}/r$ that map the semi-infinite domain $r\in\left[r_{H},\infty\right)$ to a finite interval $x\in\left[-1,1\right]$. This coordinate transformation regularizes the event horizon at $x=-1$ while preserving asymptotic flatness at $x=1$. Then the BH mass becomes
@@
[…]
@@ III Numerical method
In the compactified coordinate, the boundary conditions at horizon $x=-1$ take the following form:
@@ IV.1 The results for different $\kappa$
The comparative analysis of scalarized BH solutions across different coupling parameters reveals two key interdependent mechanisms governing their existence domains. First, the constant $\kappa$ plays a decisive role in shaping the scalarization phase space through its inverse relationship with the magnitude of the scalar charge ($Q_{s}/\sqrt{\alpha_{1}}$). This is clearly reflected in the structure of the stratified solution, where the lower $\kappa$ values correspond to upper-branch solutions with enhanced scalar charges. In addition, the quadratic coupling parameter $\alpha_{2}$ systematically suppresses scalarization across all $\kappa$ regimes. For fixed $\chi$ and $\kappa$, an increase in $\alpha_{2}$ reduces both the scalar charge and the parametric range over which viable scalarized configurations exist. Second, rotational dynamics introduces a spin-dependent suppression mechanism. At fixed values of $\kappa$ and $\alpha_{2}$, increasing the dimensionless spin $\chi$ from 00 to $0.8$ reduces the existence domain of the solution. This spin suppression becomes particularly dominant at higher $\chi$ values. To isolate the $\alpha_{2}$-$\chi$ interaction while maintaining sufficient scalarization strength for accurate measurement, we focus on case $\kappa=6$ for a further detailed analysis of the properties of the solution in the following section.
@@ IV.2 The results for $\kappa=6$
Fixing the spin $\chi=0$, Fig. 2 shows both the scalar charge and the entropy of the scalarized solution. In the no-rotating limit, the rotating scalarized solution reduces to the spherically symmetric cases, and consequently, Fig. 2 is identical to its counterpart in [63]. Each curve in the figure originates from an existence solution (right boundary point) and terminates at a critical solution (left boundary point). As shown in the right panel, the entropy of a scalarized BH exceeds that of a Schwarzschild BH with the same mass, with the exception of the coupling ratios $\alpha_{2}/\alpha_{1}=0.5$ and $0.7$ at small masses. This observation indicates that scalarized BHs are entropically favored over Schwarzschild BHs in most cases. Furthermore, as the coupling ratio $\alpha_{2}/\alpha_{1}$ increases, the mass window for the scalarized BHs shrinks. Thus, we can conclude that the additional quadratic Gauss-Bonnet term suppresses the mass window for the scalarized BHs. In the following sections, we provide a more detailed discussion of the effect of the quartic Gauss-Bonnet term on rotating scalarized BHs.
@@ IV.2.1 $\alpha_{2}/\alpha_{1}=0$
Figure 6: Metric functions $h$, $W$ and scalar fields for scalarized rotating BH solution with the same parameters as in Fig. 5. The deviations between the scalarized BH and the Kerr BH are described by $\Delta h=h-h_{Kerr}$ and $\Delta W=W-W_{Kerr}$
@@ V Conclusions
In this study, we explored how rotating BHs behave in the context of a class of sGB gravity theories that incorporate a quadratic Gauss-Bonnet term. We focus on three key factors: (1) the constant $\kappa$, which corresponds to the coupling strength between the gravitational field and scalar field; (2) the dimensionless spin parameter ($\chi$), which measures how fast a BH rotates, (3) and the coupling constant ($\alpha_{2}$), which represents additional gravitational effects in sGB gravity. Our goal is to understand how these factors influence the existence and properties of scalarized BHs. By numerically solving the field equations under stationary and axisymmetric conditions, we demonstrated that the constant $\kappa$ plays a critical role in shaping the characteristics of scalarized BHs in sGB gravity. Smaller $\kappa$ values enhance the coupling strength, leading to larger scalar charges. Moreover, we demonstrated that the scalarization of BHs is significantly suppressed by the spin parameter ($\chi$) and the quadratic coupling constant ($\alpha_{2}$). Specifically, increasing $\alpha_{2}$ narrows the domain of existence for scalarized solutions, particularly for lower-mass BHs. This suppression arises from the $\mathcal{G}^{2}$-term in the effective mass $m^{2}_{eff}$, which stabilizes general relativity solutions against tachyonic instabilities.
Furthermore, scalarized BHs are shown to be thermodynamically favored over Kerr BHs with identical mass and spin in most parameter regimes, as evidenced by their higher entropy ($S/S_{GR}$ > 1). However, this entropic preference diminished for highly spinning BHs or large $\alpha_{2}$. These results challenge the "Kerr hypothesis" and highlight the role of modified gravity in enriching the BH phase space. The predicted scalarization mass windows could be probed by future gravitational wave detectors like LISA, Taiji and TianQin, offering observational constraints on the coupling constants $\alpha_{1}$ and $\alpha_{2}$. Future work should focus on dynamical scenarios, such as BH mergers or accretion processes, to validate these predictions and refine the parameter space of sGB gravity.
@@ ACKNOWLEDGMENTS
This research is supported by the National Natural Science Foundation of China under Grant Nos.12375056, 12375048, and the Postgraduate Research &\&& Practice Innovation Program of Jiangsu Province under Grant No.KYCX223452. Some of our calculations were performed using the tensor-algebra bundle xAct [81].

### v2

# Rotating scalarized supermassive black holes

@@ Abstract
In this study, we investigate rotating black hole solutions within a scalar–Gauss–Bonnet gravity framework that incorporates a squared Gauss–Bonnet term. By employing a quadratic–exponential coupling function between the scalar field and the Gauss–Bonnet invariant, we derive both the standard General Relativity solutions and novel scalarized black hole configurations. Utilizing a pseudo-spectral method to solve the coupled field equations, we examine how black hole spin and coupling constants influence the existence and properties of these solutions. Our findings reveal that both the rotation of the black hole and the squared coupling term effectively constrain the parameter space available for scalarization. Moreover, we demonstrate that, over a wide range of parameters, scalarized black holes exhibit higher entropy than Kerr black holes of equivalent mass and spin, indicating that they are thermodynamically favored. These results significantly expand the phase space of black holes in modified gravity theories.
@@ I Introduction
Generally, scalarization is unattainable for BHs with intermediate and supermassive mass scales [31, 32, 33, 34, 35, 36]. Consquently, the prevailing expectation is that BHs in these mass ranges should be well described by the Kerr metric, even when solar-mass BHs are not. However, Ref. [63] considers a model in which the Gauss-Bonnet term is quadratically coupled to a scalar field. By studying the model in spherical symmetry, they found that unlike in the conventional framework, BHs can undergo spontaneous scalarization within a finite mass window, which may include supermassive BHs. Such deviations are particularly interesting because they offer a potential observation window through future gravitational wave detectors like LISA, Taiji, and TianQin to probe into the underlying gravitational theory [64, 65, 66, 67]. Building on the preliminary rotating solution analysis in Appendix A of Ref. [63], we present the first systematic parameter study of stationary axisymmetric scalarized black holes using pseudo-spectral methods. Our work quantifies scalarization suppression by spin and coupling constants, establishes entropy dominance over Kerr solutions, and maps the full existence domain beyond spherical symmetry.
In the present paper, we extend these investigations to rotating BHs. By employing a quadratic-exponential coupling function, we explore how the inclusion of a squared Gauss-Bonnet term modifies the spectrum of rotating BH solutions. Our study addresses two central questions: How does the presence of angular momentum affect the onset and domain of scalarization? And how do squared curvature corrections influence the stability and thermodynamic properties of these scalarized configurations compared to their Kerr counterparts? To answer these questions, we numerically solve the full set of modified field equations under stationary and axisymmetric assumptions using the Chebyshev pseudo-spectral method combined with a Newton-Raphson iterative scheme. Our analysis reveals that both the spin parameter and the additional squared coupling act to suppress the scalarization, thereby reducing the parameter space in which scalarized solutions exist. Nevertheless, in many regimes the scalarized BHs exhibit higher entropy than Kerr BHs of the same mass and spin, suggesting that they are thermodynamically favored.
@@ II.1 Action and basic equations
We consider a four-dimensional sGB gravity theory with squared curvature corrections, defined by the action [63]:
@@
where the prime $^{\prime}$" denotes the derivative with respect to scalar field $\phi$. The scalar $\psi$ acts as a Lagrange multiplier, which is not dynamical and satisfies the following equation
@@ II.2 Equations of motion and boundary behaviors
For numerical implementation of the Einstein-scalar system via the Chebyshev pseudo-spectral method, we introduce compactified radial coordinate $x=1-2r_{H}/r$ that map the semi-infinite domain $r\in\left[r_{H},\infty\right)$ to a finite interval $x\in\left[-1,1\right]$. This coordinate transformation regularizes the event horizon at x=1x=-1 while preserving asymptotic flatness at $x=1$. Then the BH mass becomes
@@
$A_{H}=2\pi r^{2}_{H}\int_{0}^{\pi}\,d\theta\sin\theta\frac{g\enspace\sqrt[]{h}}{f}\Bigg{|}_{x=-1}.$ (16)
@@ III Numerical method
In the compactified coordinate, the boundary conditions at horizon x=1x=-1 take the following form:
@@ IV.1 The results for different $\kappa$
The comparative analysis of scalarized BH solutions across different coupling parameters reveals two key interdependent mechanisms governing their existence domains. First, the constant $\kappa$ plays a decisive role in shaping the scalarization phase space through its inverse relationship with the magnitude of the scalar charge ($Q_{s}/\sqrt{\alpha_{1}}$). This is clearly reflected in the structure of the stratified solution, where the lower $\kappa$ values correspond to upper-branch solutions with enhanced scalar charges. In addition, the squared coupling parameter $\alpha_{2}$ systematically suppresses scalarization across all $\kappa$ regimes. For fixed $\chi$ and $\kappa$, an increase in $\alpha_{2}$ reduces both the scalar charge and the parametric range over which viable scalarized configurations exist. Second, rotational dynamics introduces a spin-dependent suppression mechanism. At fixed values of $\kappa$ and $\alpha_{2}$, increasing the dimensionless spin $\chi$ from 0 to $0.8$ reduces the existence domain of the solution. This spin suppression becomes particularly dominant at higher $\chi$ values. To isolate the $\alpha_{2}$-$\chi$ interaction while maintaining sufficient scalarization strength for accurate measurement, we focus on case $\kappa=6$ for a further detailed analysis of the properties of the solution in the following section.
@@ IV.2 The results for $\kappa=6$
Fixing the spin $\chi=0$, Fig. 2 shows both the scalar charge and the entropy of the scalarized solution. In the no-rotating limit, the rotating scalarized solution reduces to the spherically symmetric cases, and consequently, Fig. 2 is identical to its counterpart in [63]. Each curve in the figure originates from an existence solution (right boundary point) and terminates at a critical solution (left boundary point). As shown in the right panel, the entropy of a scalarized BH exceeds that of a Schwarzschild BH with the same mass, with the exception of the coupling ratios $\alpha_{2}/\alpha_{1}=0.5$ and $0.7$ at small masses. This observation indicates that scalarized BHs are entropically favored over Schwarzschild BHs in most cases. Furthermore, as the coupling ratio $\alpha_{2}/\alpha_{1}$ increases, the mass window for the scalarized BHs shrinks. Thus, we can conclude that the additional squared Gauss-Bonnet term suppresses the mass window for the scalarized BHs. In the following sections, we provide a more detailed discussion of the effect of the quartic Gauss-Bonnet term on rotating scalarized BHs.
@@ IV.2.1 $\alpha_{2}/\alpha_{1}=0$
Figure 6: Metric functions $h$, $W$ and scalar fields for scalarized rotating BH solution with the same parameters as in Fig. 5. The deviations between the scalarized BH and the Kerr BH are described by $\Delta h=h-h_{Kerr}$ and $\Delta W=W-W_{Kerr}$ .
@@ V Conclusions
In this study, we explored how rotating BHs behave in the context of a class of sGB gravity theories that incorporate a squared Gauss-Bonnet term. We focus on three key factors: (1) the constant $\kappa$, which corresponds to the coupling strength between the gravitational field and scalar field; (2) the dimensionless spin parameter ($\chi$), which measures how fast a BH rotates, (3) and the coupling constant ($\alpha_{2}$), which represents additional gravitational effects in sGB gravity. Our goal is to understand how these factors influence the existence and properties of scalarized BHs. By numerically solving the field equations under stationary and axisymmetric conditions, we demonstrated that the constant $\kappa$ plays a critical role in shaping the characteristics of scalarized BHs in sGB gravity. Smaller $\kappa$ values enhance the coupling strength, leading to larger scalar charges. Moreover, we demonstrated that the scalarization of BHs is significantly suppressed by the spin parameter ($\chi$) and the squared coupling constant ($\alpha_{2}$). Specifically, increasing $\alpha_{2}$ narrows the domain of existence for scalarized solutions, particularly for lower-mass BHs. This suppression arises from the $\mathcal{G}^{2}$-term in the effective mass $m^{2}_{eff}$, which stabilizes general relativity solutions against tachyonic instabilities.
Furthermore, scalarized BHs are shown to be thermodynamically favored over Kerr BHs with identical mass and spin in most parameter regimes, as evidenced by their higher entropy ($S/S_{GR}$ > 1). However, this entropic preference diminished for highly spinning BHs or large $\alpha_{2}$. These results challenge the "Kerr hypothesis" and highlight the role of modified gravity in enriching the BH phase space. The predicted scalarization mass windows could be probed by future gravitational wave detectors like LISA, Taiji and TianQin, offering observational constraints on the coupling constants $\alpha_{1}$ and $\alpha_{2}$. It should be noted that the rotating scalarized BHs investigated in this work fall within the curvature-induced scalarization framework. In contrast, spin-induced scalarization, which represents a distinct mechanism and has garnered significant interest in recent studies [68, 69, 70, 71], remains an intriguing direction for future exploration. Future work should focus on dynamical scenarios, such as BH mergers or accretion processes, to validate these predictions and refine the parameter space of sGB gravity.
@@ ACKNOWLEDGMENTS
This research is supported by the National Natural Science Foundation of China under Grant Nos.12375056, 12375048, and the Postgraduate Research &\& Practice Innovation Program of Jiangsu Province under Grant No.KYCX24_3712. Some of our calculations were performed using the tensor-algebra bundle xAct [81].